How many moles of FeCl2 are produced when 16 grams of Fe reacts with HCl?

1
Expert's answer
2022-04-06T08:56:04-0400

Fe + 2HCl = FeCl2 + H2

According to the equation, n (FeCl2) = 1/2 x n (HCl)

n = m / M

M (Fe) = 55.9 g/mol

n (Fe) = 16 / 55.9 = 0.29 mol

n (FeCl2) = 1/2 x n (HCl) = 1/2 x 0.29 = 0.14 mol